OpenStudy (reginabusch):

The graph of f(x) = 2x + 4 shifts six units to the right when it is replaced with the graph of f(x) = 2x - k. What is the value of k? 2 4 6 10

OpenStudy (welshfella):

first lets find where the graph f(x) = 2x + 4 cuts the x axis 2x + 4 = 0 where it cuts x-axis so 2x + 4 = 0 2x = -4 x = -2

OpenStudy (reginabusch):

I figured it out, thank you so much!

OpenStudy (welshfella):

So the new line moves 6 to the right therefore it cuts the xx-axis at x = 4 so we have 0 = 2(4) - k solve for k

OpenStudy (welshfella):

ok but none of the choices are correct k = 8

OpenStudy (reginabusch):

hmm I got 6 and it was correct. I guess theres a typo in the question. Thank you anyways

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

if k = 6 then the lines moves 5 units to the right